### v1

# Towards a more realistic evaluation of machine learning models for bearing fault diagnosis

Abstract
Reliable detection of bearing faults is essential for maintaining the safety and operational efficiency of rotating machinery. While recent advances in machine learning (ML), particularly deep learning, have shown strong performance in controlled settings, many studies fail to generalize to real-world applications due to methodological flaws, most notably data leakage. This paper investigates the issue of data leakage in vibration-based bearing fault diagnosis and its impact on model evaluation. We demonstrate that common dataset partitioning strategies, such as segment-wise and condition-wise splits, introduce spurious correlations that inflate performance metrics. To address this, we propose a rigorous, leakage-free evaluation methodology centered on bearing-wise data partitioning, ensuring no overlap between the physical components used for training and testing. Additionally, we reformulate the classification task as a multi-label problem, enabling the detection of co-occurring fault types and the use of prevalence-independent metrics such as Macro AUROC. Beyond preventing leakage, we also examine the effect of dataset diversity on generalization, showing that the number of unique training bearings is a decisive factor for achieving robust performance. We evaluate our methodology on three widely adopted datasets: CWRU, Paderborn University (PU), and University of Ottawa (UORED-VAFCLS). This study highlights the importance of leakage-aware evaluation protocols and provides practical guidelines for dataset partitioning, model selection, and validation, fostering the development of more trustworthy ML systems for industrial fault diagnosis applications.
keywords:
[…]
1 Introduction
The field of fault diagnosis for rotating machinery, particularly rolling bearings, has seen increased attention due to its critical role in various industries and the demand for efficient operations [1]. Early and accurate detection of bearing failures can significantly reduce unexpected machine downtime and improve maintenance schedules, avoiding financial losses and safety risks. Machine Learning (ML) approaches, including deep learning architectures, coupled with wireless sensor technologies, have enabled health monitoring and failure prediction at scale.
Despite the significant advancements offered by ML, its application requires a careful methodology to ensure models generalize reliably to real-world scenarios. A critical methodological pitfall in ML-based science is data leakage [2]. Data leakage is defined as a spurious relationship between independent variables and the target variable, arising from flaws in data collection, sampling, or preprocessing [3]. Such an artifact, not present in the true data distribution, typically leads to overoptimistic estimates of model performance. This phenomenon is a major source of error in ML applications, often causing published models to fail when deployed in practical settings.
The problem of data leakage is not unique to bearing fault diagnosis; it is a widespread failure mode in ML-based science, impacting at least 294 papers across 17 scientific fields [3]. For instance, in medicine, improper handling of patient data can lead to leakage if samples from the same patient are used in both training and test sets. Cases have been reported where models included features that were effectively proxies for the outcome, such as the use of anti-hypertensive drugs to predict hypertension or antibiotics to predict sepsis, leading to artificially inflated performance.
Our observations indicate that the problem of data leakage is prevalent in the area of bearing fault diagnosis. Many studies have failed to properly partition their datasets, leading to information leakage and consequently, over-optimistic performance estimates that do not hold in real-world scenarios. For example, studies assigning waveform recordings from the same bearing to both training and test partitions have reported inflated performance. Early work by [4] highlighted this “similarity bias” in ML research using vibration data, finding that nearly all reviewed studies, including 40 out of 41 using the widely adopted Case Western Reserve University (CWRU) dataset, employed experimental designs susceptible to this bias. Recent publications continue to show overestimated results due to similarity bias. Comprehensive reviews in vibration-based intelligent fault diagnosis have also largely overlooked this crucial issue.
Previous efforts have attempted to mitigate data leakage in bearing fault diagnosis. Hendriks et al. [5], for instance, proposed a dataset split for CWRU based on fault size rather than motor load, which significantly reduced the inflated classification accuracy observed in earlier studies. Similarly, [6] explored alternative data splits, demonstrating that assigning vibration data from a given bearing across training and evaluation splits leads to over-optimistic performance estimates. Their work showed notable performance drops (e.g., Naive Bayes accuracy decreased from 85.8% to 69.5% when moving from a random split with leakage to a bearing-wise split) when leakage was addressed. Despite these valuable contributions, a common limitation remained on the CWRU dataset: their multi-class formulations can not entirely eliminate leakage from the healthy class, as this dataset contains only one healthy bearing configuration. Additionally, since the CWRU dataset contains only faulty bearings with a single fault, their multi-class approach cannot account for scenarios where multiple fault types might occur simultaneously.
Given these challenges, this paper aims to further advance the reliable development and deployment of ML models for bearing fault diagnosis. We propose a novel methodology that rigorously addresses data leakage and class imbalance, particularly for datasets with limited healthy bearing data. Our approach formulates the problem as a binary multi-label classification for each sensor location (e.g., drive end and fan end), enabling the detection of the presence or absence of each fault type (e.g., inner, outer, ball). This formulation enables a more precise evaluation by using prevalence-independent metrics such as the Receiver Operating Characteristic (ROC) curve and the Area Under the Receiver Operating Characteristic curve (AUROC). Furthermore, it provides a more realistic representation of real-world conditions where multiple fault types could coexist. Our contributions include:
  • Demonstrating the significant impact of data leakage on classification accuracy across different train/test splits, revealing that models that appear to perform well with leakage may exhibit a substantial drop in performance (e.g., over 40% accuracy drop in some cases) when evaluated without it.
  • Providing a systematic methodology for creating and using vibration-based datasets in ML experiments that strives to prevent data leakage and suggesting an accompanying hyperparameter tuning process that adheres to the same bias minimization principle.
  • Proposing a multilabel problem formulation, which enables a more precise evaluation by using prevalence-independent metrics such as the ROC curve and the AUROC and provides a more realistic representation of real-world conditions where multiple fault types could coexist.
  • Applying the proposed methodology on widely-used vibration datasets, such as CWRU, Paderborn University (PU), and University of Ottawa (UORED-VAFCLS) datasets. Our results reveal the significant impact of bearing diversity on model generalization and demonstrate that the optimal choice between deep and shallow learning models is highly dataset-dependent.
With these contributions, this work aims to foster the development of more robust and trustworthy ML models for bearing fault diagnosis, ensuring their performance more closely reflects their capabilities in real-world industrial settings. The source code for this paper can be found at https://github.com/gama-ufsc/bearing-data-leakage.
2 Background
2.1 Basic concepts on supervised learning
Supervised machine learning is a paradigm centered on learning a mapping from inputs to outputs based on a set of labeled examples. The fundamental goal is to approximate an unknown underlying function that dictates the relationship between the observed data and their corresponding labels.
Mathematically, we consider an input space $\mathcal{X}$ and an output space $\mathcal{Y}$. The relationship between them is governed by a true, but unknown, joint probability distribution $P(X,Y)$, where $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$. We are not given access to $P(X,Y)$ directly. Instead, we are provided with a finite set of observations, known as the training set, $D_{train}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where each sample $(x_{i},y_{i})$ is assumed to be an independent and identically distributed (i.i.d.) draw from $P(X,Y)$.
Each input $x_{i}$ is typically a feature vector, $x_{i}\in\mathbb{R}^{d}$, representing a set of $d$ measurable properties of the phenomenon being observed. The output $y_{i}$ is the corresponding label or target value. The task is to select a model from a hypothesis space $\mathcal{H}$, which is a family of functions $f:\mathcal{X}\to\mathcal{Y}$. A specific model is defined by a set of parameters, $\theta$, denoted as $f_{\theta}$.
The learning process consists of finding the optimal parameters $\theta^{*}$ that enable the model to make accurate predictions. To achieve this, we first define a loss function, $\mathcal{L}(f_{\theta}(x),y)$, which quantifies the penalty or error for predicting $f_{\theta}(x)$ when the true label is $y$. The ultimate objective is to minimize the true risk or expected loss over the entire data distribution
$R(f_{\theta})=\mathbb{E}_{(x,y)\sim P(X,Y)}[\mathcal{L}(f_{\theta}(x),y)].$
Since $P(X,Y)$ is unknown, the true risk cannot be calculated directly. Therefore, we approximate it using the empirical risk on the training set
$R_{emp}(f_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i}).$
The training process then becomes an optimization problem focused on finding the parameters $\theta^{*}$ that minimize this empirical risk
$\theta^{*}=\arg\min_{\theta}R_{emp}(f_{\theta}).$
The model’s parameters, $\theta^{*}$, are optimized exclusively on the training set, $D_{train}$. Since the model is ultimately a function of the training set, the model’s performance on that same data is not a reliable indicator of its actual predictive power; it is an inherently biased and optimistic measure. What truly matters is the model’s generalization performance—how well it performs on new, unseen data from the same underlying distribution, $P(X,Y)$. To perform this evaluation, we use a disjoint test set, $D_{test}$, which is a collection of i.i.d. samples from $P(X,Y)$ kept completely separate during the entire training and model selection process.
However, in addition to the learnable parameters $\theta$, most models are also characterized by a set of hyperparameters, $\lambda$, which are not optimized during training but rather define the model’s architecture or the learning algorithm’s behavior (e.g., learning rate, regularization strength). The process of selecting the optimal configuration of these hyperparameters is known as model selection. Using the test set, $D_{test}$, to guide this selection process is methodologically unsound, as it would mean that information from the test set has leaked into the model configuration, violating the i.i.d. assumption. Consequently, $D_{test}$ would no longer provide an unbiased estimate of the final model’s generalization performance.
To address this, the dataset $D$ is typically partitioned into three disjoint subsets: a training set ($D_{train}$), a validation set ($D_{val}$), and a test set ($D_{test}$). The hyperparameter optimization process proceeds as follows: for each candidate hyperparameter configuration $\lambda$, a model is trained on $D_{train}$ to find the optimal parameters $\theta^{*}(\lambda)$. The performance of this trained model is then evaluated on the validation set, yielding a validation risk. The hyperparameter configuration $\lambda^{*}$ that results in the lowest validation risk is selected as the optimal one:
$\lambda^{*}=\arg\min_{\lambda}\left(\frac{1}{|D_{val}|}\sum_{(x,y)\in D_{val}}\mathcal{L}(f_{\theta^{*}(\lambda)}(x),y)\right).$
Once the optimal hyperparameters $\lambda^{*}$ have been identified, the final model is trained and its generalization performance is reported based on a single, final evaluation on the held-out test set, $D_{test}$. The performance on this held-out data gives us an unbiased estimate of the true risk, $R(f_{\theta^{*}(\lambda^{*})})$. In essence, it is our best proxy for how the model will behave in the wild. This practice is crucial for diagnosing overfitting, a common pitfall where a model appears highly accurate on the data it was trained on but fails to generalize to new examples.
While the assumption that data samples are independent and identically distributed is fundamental to supervised learning, it is frequently violated in real-world machine learning applications, potentially leading to overly optimistic performance estimates if not properly addressed [7]. A common scenario where this assumption breaks down is when the underlying data generation process naturally produces groups of dependent samples, such as multiple medical records from the same patient [3]. Addressing these violations is critical for achieving robust and reliable model performance, preventing these overly optimistic estimates. The overarching strategy involves group cross-validation (CV) [8] [^1] An implementation of cross-validation applied to grouped data is available on https://scikit-learn.org/stable/modules/cross_validation.html#cross-validation-iterators-for-grouped-data., where data is split in a manner that respects these inherent correlation structures rather than through purely random partitioning, ensuring that training and evaluation data remain genuinely independent.
3 Data Leakage
3.1 Data splitting and Data Leakage
Data leakage is a major source of error in machine learning applications [9], and often the reason why published models fail to generalize to real-world data [10]. As defined by [3], it refers to spurious correlations between input and target variables arising from flaws in data collection, sampling, or preprocessing. These artificial relationships, absent in the true data distribution, typically yield overly optimistic performance estimates during development but poor generalization to unseen data.
As mentioned in Section 2, there are many methodologic flaws that violate the i.i.d. assumption, which ultimately results in data leakage. This may happen through model or feature selection before data partitioning, during preprocessing, improper handling of test data and splitting methods. In sequence, examples for the mentioned cases will be discussed.
A common flaw that introduces data leakage during preprocessing is computing statistics such as means or ranges for scaling, or imputing missing values, using the entire dataset instead of restricting calculations to the training portion. Similarly, performing model or feature selection before data partitioning allows information from the test set to influence model configuration [10]. To prevent these issues, all preprocessing and feature engineering steps must be fitted exclusively on training data.
Another source of data leakage comes from improper handling of test data. Using the same test set to evaluate multiple models can inadvertently inform model selection, leading to overfitting. Additionally, applying data augmentation before dividing the dataset can cause augmented information from the test set to seep into the training data, compromising the model’s ability to generalize.
In time-series applications, random partitioning without preserving temporal order can allow future information to leak into the training process, producing inflated performance metrics. Even in non-temporal datasets, experimental designs may introduce dependencies or duplicate samples that, if split incorrectly, create information overlap between train and test sets.
Key to avoiding these pitfalls is strict maintenance of train–test separation throughout the entire pipeline. This means ensuring that no information from the test set influences preprocessing, feature selection, hyperparameter tuning, or model training [11]. In practice, this involves grouping dependent samples (e.g., patient-level records or bearing-level measurements) within the same fold, and using specialized cross-validation schemes for structured data—such as blocked CV for time-series [7]—to prevent “look-ahead bias” and maintain a valid evaluation protocol.
3.1.1 Data Leakage in Bearing Fault Diagnosis
In the evaluation of bearing fault diagnosis models, data leakage represents a significant pitfall that can lead to overly optimistic performance estimates. We identify and categorize two common but flawed intra-bearing splitting protocols that lead to various forms of leakage: segmentation-level and bearing-level leakage.
Segmentation-Level Leakage occurs when non-overlapping segments from the same time-series signal (e.g., from a continuous experimental run) are split between training and test sets. Although the segments do not overlap in time, the underlying signal is temporally and physically coherent, allowing the model to learn time-specific or signal-specific artifacts rather than generalizable fault features. In bearing fault diagnosis, this form of leakage is especially common and is discussed in detail in [12].
Bearing-Level Leakage arises when data from the same physical bearing is distributed across both training and test sets. This may occur through common random splitting procedures or, more specifically, through approaches such as condition-wise and repetition-wise splits, defined as follows.
A condition-wise split happens when signals with different machine conditions are divided between training and testing. A condition is usually represented by the combination of machine configurations, such as the load and rotation speed, for instance. It can also be represented by a severity level in a given bearing, as in the UORED-VAFCLS dataset, where two fault severity levels are available for every bearing. This type of data splitting is commonly used in bearing fault diagnosis, as it is generally understood that a realistic evaluation methodology requires diverse machine configurations to be separated between training and testing. While some may argue that such variation justifies this split, it is conceivable that some intrinsic signature of the bearing remains present in both sets, allowing the model to exploit identity-specific features rather than learning robust patterns.
An even more severe form of data leakage arises from repetition-wise splits (also referred to as run-to-run in [13]), where signals acquired under the same configuration are divided between training and testing. Since signals captured under identical setups typically share similar characteristics, this setup enables models to memorize the signals rather than learn to diagnose faults accurately. In Section 6.2, we show that in the PU dataset, all splits with bearing-level leakage produce overoptimistic results, with repetition- and segment-wise splits reaching almost 100% accuracy.
All these cases result in evaluations that are fundamentally closer to measuring memorization (e.g. the training performance) rather than generalization. To preclude such pitfalls, we advocate for a strict bearing-wise split, in which the physical bearings used for training and testing are mutually exclusive. This approach ensures the model is assessed on its ability to generalize to unseen components, preventing it from leveraging identity-specific artifacts [^2] If samples from the same bearing but with different labels (such as when a healthy bearing is subject to accelerated lifetime tests and develops a fault) are separated into training and test sets, this would not characterize bearing-level leakage, since the spurious correlation between them would not favor the model’s decision towards the correct class. Instead, it would constitute a more challenging evaluation process, which tests whether the model has learned the correct feature representation rather than relying on bearing signatures.. Ultimately, this promotes the development of models that identify universal fault characteristics, a goal best supported by datasets that include multiple physical bearings per fault category. It is important to note that even with bearing-wise splits, leakage may still occur due to spurious correlations learned from poor data collection practices. For example, if all fault signals in a dataset contain additional interference absent from healthy signals, the model might rely on that interference for classification, thus enabling memorization.
3.2 A Toy Example
To demonstrate the performance inflation caused by a non-bearing-wise data split, we constructed a synthetic binary fault detection experiment, in which the task is to classify each sample as either faulty or healthy. We simulate a feature space for a dataset of $B=48$ distinct bearings, which could be obtained, for example, as the output of a deep learning model’s feature extractor. This simulated space contains two feature types: 3 fault-predictive features and 48 bearing-identity features, totalizing 51 features. Each identity feature is unique to a specific bearing, representing a spurious correlation unrelated to the fault condition. All features were generated by adding zero-mean Gaussian noise with unit variance to a constant base value: $a_{f}=1.5$ for fault-predictive features and $a_{b}=8$ for bearing-identity features. All experiments were designed to maintain a fixed number of samples per bearing, using a base value of 40 samples each. The dataset comprised 24 healthy bearings and 24 faulty bearings, ensuring a consistently balanced class distribution.
First, we established a theoretical performance ceiling for the classifier, as shown in A. In summary, computing the mean of the 3 fault predictive features and applying a threshold of $a_{f}/2$ gives the maximum achievable accuracy of $90.30\%$. We then evaluated models with two distinct test set configurations: the first consisted of samples generated from the same bearings used in the training set, while the second comprised samples generated from different bearings.
Using Logistic Regression and Decision Tree classifiers to assess the impact of model capacity, our results (Figure 1) show that test sets containing samples from bearings included in the training set yield overly optimistic performance metrics for both classifiers, even exceeding the theoretical maximum classification accuracy. Due to data leakage, the results misleadingly suggest that Logistic Regression is the superior model; however, under the valid test set, the Decision Tree actually demonstrates better performance when 4 or more training bearings are used. This highlights that results influenced by data leakage are unreliable for informed decision-making. Additionally, increasing the number of bearings in training improves performance on the leakage-free test set, while it also mitigates the impact of leakage on both models, suggesting it learns to prioritize the correct predictive features. These findings underscore the critical importance of a rigorous validation methodology to prevent misinterpretation of a model’s true generalization capabilities.
Figure 1: Comparison of Decision Tree (DT) and Logistic Regression (LR) accuracy across varying numbers of training bearings, evaluated under two conditions: a leakage-free test (Valid) set and a test set with data leakage (Leakage).
3.3 Related works
Several studies in the literature have highlighted the issue of data leakage in bearing fault diagnosis, including [13], [5], [6], and [12]. These works propose new data-splitting strategies that primarily avoid segmentation-level leakage, but still allow bearing-wise leakage to persist (with the exception of [13]).
Hendriks et al. [5] proposed a bearing-wise splitting strategy on the CWRU dataset. Their methodology closely resembles ours, with one critical difference: healthy signals were included in both training and testing sets without applying consistent separation criteria, resulting in data leakage from the healthy bearings.
Abburi et al. [6] also proposed an alternative split strategy aimed at reducing bearing-level information leakage on the CWRU dataset. The authors employed traditional machine learning models in a multiclass classification setting with three fault types (inner race, outer race, ball) and the healthy condition. Their results showed that the bearing-wise split consistently led to worse performance across all metrics (accuracy, precision, recall, and macro F1-score) when compared to a random split that included bearing information leakage. However, multiclass formulations such as this are unable to fully prevent leakage from the healthy class in the CWRU—a limitation we address through our proposed multi-label framework. To address the issues on [5, 6], we introduce a revised split strategy for CWRU, detailed in Section 4.2.3.
Another notable contribution is the study by [13], which proposed a bearing-wise split under a multiclass classification setting, referred to as a “part-to-part” approach. Using the KAt and CMTH datasets, the authors demonstrated a high correlation between signals originating from the same bearing—violating the i.i.d. assumption—and showed the substantial drop in accuracy when shifting from a condition-wise to a bearing-wise split, revealing the effects of a bearing-level leakage.
A comprehensive investigation into dataset biases and evaluation protocols was conducted by [12], who proposed evaluation methodologies for widely used datasets such as CWRU, PU, MFPT, IMS, and UOC. Their approach focused on mitigating segmentation-level leakage, advocating for condition-wise splits across all datasets under a multiclass framework, with F1-macro reported as an auxiliary metric. However, the results remained overoptimistic due to residual bearing-level leakage.
It is worth mentioning that a few studies using the PU dataset have apparently adopted strict bearing-wise splits and reported strong results, such as [14] and [15]. This approach was also employed in the original PU paper [16], which reported a 98.3% multiclass accuracy. Although these works provide clear instructions on their proposed splitting strategies, we were unable to reproduce the same results in our own experiments.
4 Methodology
In this section, we first introduce our general methodology, including a bearing-wise split example using a generic dataset, the evaluation metric, and the hyperparameter tuning and model evaluation protocol. We then apply and specialize this methodology to three public bearing diagnosis datasets: University of Ottawa (UORED-VAFCLS), Paderborn University (PU), and Case Western Reserve University (CWRU). Finally, we describe the features, deep learning architectures, and data augmentation techniques used in our experiments.
4.1 General Methodology
The problem of detection and diagnosis of bearing faults can be formulated in many different ways, which can profoundly impact what a model can predict and how it is evaluated. Based on our findings regarding data leakage, we propose a methodology founded on a bearing-wise data splitting with the adoption of a multi-label classification framework to accommodate the potential for co-occurring fault modes and mitigate the effect of class imbalance. In a broader context, our proposed multi-label approach to fault detection and diagnosis is that of constructing specific detectors for each possible type of fault. Then, conventional fault detection amounts to simply verifying if any of the specific detectors gives a positive output, while fault diagnosis amounts to retrieving which specific detectors give a positive output. Additionally, this approach facilitates the use of sophisticated binary classification metrics, such as the Area Under the ROC Curve (AUC-ROC), which provide a threshold-independent and more nuanced evaluation of diagnostic performance compared to standard accuracy scores. A single summary metric, the Macro AUROC, is obtained by averaging the AUROCs of all individual detectors.
4.1.1 Data Splitting
Central to our methodology is a strict, bearing-level data partitioning strategy designed to prevent data leakage and ensure a valid assessment of model generalization. To illustrate this principle, we define a generic dataset construct, specified in Figure 2. This dataset comprises $B=15$ unique physical bearings, each characterized by one of three health states: healthy, inner race fault, or outer race fault.
Figure 2: Exemplary bearing-level data partitioning for the generic dataset. The training set (green) and test set (blue) are disjoint at the bearing level, with a 3:2 allocation of bearings per health state.
Under our multi-label framework, these states are represented by binary vectors, where a healthy bearing is encoded as [0,0], an inner race fault as [1,0], and an outer race fault as [0,1]. The partitioning of this dataset, which is detailed in Figure 3, adheres to a 3:2 train-to-test ratio applied at the bearing level. Specifically, for each health state, three distinct bearings are allocated to the training set, while the remaining two are reserved for the test set. This ensures that no data from a single physical bearing appears in both the training and test partitions, thereby creating a realistic scenario for evaluating performance on unseen components.
Figure 3: Specification of the generic bearing fault dataset, comprising 15 unique bearings, two fault modes (inner, outer), and two distinct acquisition configurations per bearing.
4.1.2 Evaluation metric
Although multiclass accuracy is a common metric in the cited literature, it has significant limitations for bearing fault diagnosis. Firstly, it is sensitive to class imbalance, a prevalent issue in fault diagnosis datasets. While balanced accuracy can mitigate this by weighting classes, it introduces another problem: the fault frequencies in benchmark datasets rarely mirror real-world conditions, where healthy states dominate until a fault develops. Secondly, the metric’s formulation assumes mutually exclusive classes, which inherently prevents the diagnosis of co-occurring faults. Creating additional classes for every fault combination is often impractical, as these events are too rare in datasets for a model to learn effectively. Finally, multiclass accuracy lacks a tunable decision threshold, making it impossible to optimize for specific operational needs, such as minimizing the False Negative Rate to avoid missed detections.
A key advantage of the proposed multi-label framework is its compatibility with evaluation metrics that are robust to class imbalance. By treating the detection of each fault type as an independent binary classification problem, we can employ fine-grained, prevalence-independent metrics. The primary metric selected for this work is the Area Under the Receiver Operating Characteristic curve (AUROC). The AUROC provides a threshold-independent measure of a classifier’s ability to distinguish between positive and negative classes, making it particularly suitable for bearing fault datasets where healthy samples often vastly outnumber faulty ones.
To obtain a single, comprehensive measure of overall model performance across all $F$ fault types, we compute the macro-average of the individual AUROC scores. This aggregate metric is calculated as
$\text{Macro AUROC}=\frac{1}{F}\cdot\sum_{i=1}^{F}\text{AUROC}_{i}.$
4.1.3 Hyperparameter Tuning and Model evaluation
For hyperparameter optimization (also called model selection) and performance evaluation, we adopt the Double Cross-Validation Method (CVM-CV), described in [17]. This protocol consists of applying the cross-validation method for hyperparameter optimization (CVM) and then reevaluating it on different train-test splits (CV) for final performance estimation for the single, selected, best model. Note that using only CVM is well-known to overestimate performance since it returns the maximum performance achieved across several hyperparameter configurations. This bias can be reduced by reevaluating only the selected hyperparameter configuration on different train-test splits.
Our implementation follows a two-stage process:
  • Hyperparameter Optimization (CVM): An inner cross-validation loop is employed exclusively for identifying the optimal hyperparameter set for a given model. To accommodate the specific structures of the public datasets, this stage was adapted: for the PU and OU datasets, a 5-run random train-test split was used, while for the more constrained CWRU dataset, a 3-fold partitioning was applied. The hyperparameter configuration yielding the highest average performance in this inner loop was selected for the next stage.
  • Performance Estimation (CV): The model, using the selected hyperparameters, is retrained and evaluated across 100 distinct, randomly generated train-test splits. Although some test data in this stage may have been seen during hyperparameter tuning, the large number of disjoint splits dilutes the influence of any specific instance. The final reported performance is the average metric (e.g., Macro AUROC) over these 100 runs, providing a more stable and representative estimate of the model’s generalization ability.
4.2 Specific details for each dataset
4.2.1 University of Ottawa (UORED-VAFCLS)
The University of Ottawa Rolling-element Dataset – Vibration and Acoustic Faults under Constant Load and Speed conditions (UORED-VAFCLS) [18] provides a contemporary benchmark for fault diagnosis methodologies, offering multi-modal data streams including acoustic, vibration, and temperature signals. The dataset encompasses four distinct fault modes: inner race, outer race, ball, and cage. For each fault category, five unique physical bearings were tested, resulting in a total of 20 distinct components. All acquisitions were conducted under a single, fixed operating condition (500N load, 1750 RPM) with signals recorded for 10 seconds at a 42 kHz sampling rate.
A notable characteristic of the UORED-VAFCLS dataset is its hierarchical structure: each of the 20 physical bearings was recorded across three progressive health states: 1) normal operation, 2) weak fault severity, and 3) strong fault severity. This design results in a total of 60 discrete time-series signals. The composition of the dataset is summarized in Table 1.
Table 1: Bearing-level structure of the University of Ottawa (UORED-VAFCLS) dataset, detailing the 20 unique bearings across four fault categories.
Bearing ID Health state Bearing ID Health state Bearing ID Health state Bearing ID Health state
1 Healthy 6 Healthy 11 Healthy 16 Healthy
1 Inner-1 6 Outer-1 11 Ball-1 16 Cage-1
1 Inner-2 6 Outer-2 11 Ball-2 16 Cage-2
2 Healthy 7 Healthy 12 Healthy 17 Healthy
2 Inner-1 7 Outer-1 12 Ball-1 17 Cage-1
2 Inner-2 7 Outer-2 12 Ball-2 17 Cage-2
3 Healthy 8 Healthy 13 Healthy 18 Healthy
3 Inner-1 8 Outer-1 13 Ball-1 18 Cage-1
3 Inner-2 8 Outer-2 13 Ball-2 18 Cage-2
4 Healthy 9 Healthy 14 Healthy 19 Healthy
4 Inner-1 9 Outer-1 14 Ball-1 19 Cage-1
4 Inner-2 9 Outer-2 14 Ball-2 19 Cage-2
5 Healthy 10 Healthy 15 Healthy 20 Healthy
5 Inner-1 10 Outer-1 15 Ball-1 20 Cage-1
5 Inner-2 10 Outer-2 15 Ball-2 20 Cage-2
This multi-state-per-bearing structure does not compromise the integrity of our proposed methodology. On the contrary, it underscores the necessity of bearing-level partitioning. By assigning all signals from a single physical component exclusively to either the training or the test set, we rigorously prevent data leakage and ensure a valid assessment of the model’s ability to generalize to entirely unseen hardware.
To implement our CVM-CV protocol, we systematically partitioned the dataset. For each of the four fault modes, the five available bearings were split into a 3:2 train-test ratio. The number of unique ways to select three of the five bearings for the training set is equal to 10. As the selection for each fault mode is independent, the total combinatorial space of unique splits is 104. From this space, we instantiated 105 distinct splits for our experiment.
As illustrated in Figure 4, these splits were strictly segregated:
  • Hyperparameter Tuning (CVM): The first 5 unique splits were used exclusively for the inner cross-validation loop to perform model selection.
  • Performance Estimation (CV): The subsequent 100 disjoint splits were reserved for the outer loop to evaluate the performance of the selected model.
This two-tiered approach guarantees that the data combinations used for performance estimation were entirely unseen during the hyperparameter tuning process, thereby yielding a more robust measure of model generalization.
Figure 4: Schematic of the Double Cross-Validation (CVM-CV) protocol applied to the UORED-VAFCLS dataset. A distinct set of 5 bearing-level splits is used for hyperparameter tuning, while a separate set of 100 splits is used for final performance evaluation.
4.2.2 Paderborn University (PU) Dataset
The Paderborn University (PU) bearing dataset [16] represents a complex and widely-used benchmark, distinguished by its inclusion of bearings from multiple manufacturers and two distinct fault origination paradigms: artificial damage and natural degradation from accelerated lifetime tests. The dataset contains 6 healthy bearings, 12 with artificially induced faults (inner/outer race), and 14 with faults developed during operation (inner race, outer race, and combined inner/outer race). Data was captured under four discrete operating conditions, varying rotational speed, load torque, and radial force, with vibration signals recorded at a 64 kHz sampling rate.
Crucially, prior work by [16] demonstrated that a significant domain shift exists between artificially damaged and naturally degraded bearings, leading to poor generalization when training on the former and testing on the latter. To circumvent this issue and ensure the practical relevance of our findings, our investigation exclusively utilizes the subset of bearings with naturally occurring faults from accelerated lifetime tests, along with the healthy reference bearings. Our analysis incorporates components K006 and KI17, thereby expanding upon the set used in the original benchmark study, which are shown in Table 2. Although bearings with combined faults are in principle compatible with our multi-label methodology, their limited representation in the dataset hinders robust learning and evaluation. In fact, only three such bearings are available (IDs KB23, KB24, and KB27). Therefore, we exclude these samples from both training and testing to prevent biased or unreliable conclusions.
Table 2: IDs of all healthy and naturally damaged bearings from the PU dataset utilized in this study. The selection expands upon the original benchmark set from [16] (highlighted in yellow) by incorporating additional available components (highlighted in green).
Healthy
Outer ring
damage
Inner ring
damage
K001 KA04 KI04
K002 KA15 KI14
K003 KA16 KI16
K004 KA22 KI18
K005 KA30 KI21
K006 - KI17
Our partitioning strategy was tailored to the specific composition of this curated subset. To accommodate the varying number of available bearings along the three classes, a differentiated partitioning scheme was adopted for the CVM-CV protocol. For the healthy and inner race fault categories, each of which contains six bearings, a 4:2 train-test split was implemented. For the outer race fault category, with five available bearings, a 3:2 split was used. As depicted in Figure 5, the splits designated for hyperparameter tuning and final performance evaluation were kept entirely separate to maintain the integrity of the validation process.
Figure 5: Schematic of the Double Cross-Validation (CVM-CV) protocol applied to the PU dataset.
4.2.3 CWRU bearing fault dataset
The Case Western Reserve University (CWRU) bearing fault dataset [19] is a collection of experiments that involved a single pair of healthy bearings and several artificially created faulty bearings. The faults were created through electro-discharge machining, introducing point faults with diameters of 7, 14, 21, and 28 mils in the inner race, outer race, and rolling element separately. For the outer race faults, the experiments considered faults located at three different positions relative to the load zone. The healthy and faulty bearings were reinstalled at both the drive end (DE) and fan end (FE) locations (where each configuration comprises either two healthy bearings or one healthy bearing and one faulty bearing), and data were collected synchronously, with one accelerometer at each location. For each configuration, experiments were made using four operational motor load conditions ranging from 0 (no load) to 3 horsepower (HP). In most cases, the experiments used a 12 kHz sampling rate, while some used 48 kHz. All the experiments consist of signals that are approximately 10 seconds long.
Following a similar approach to [5], the configurations considered in this paper used a load varying from 0 to 3 HP and fault sizes of 7, 14 and 21 mils. Measurements with a sampling rate of 12 kHz were used, except for the healthy bearing experiments that only had a sampling rate of 48 kHz available and were resampled to 12 kHz. Considering the three different fault positions at the outer race fault experiments, the “Centered @6:00” experiments were primarily used whenever possible. If the former did not exist, the “Orthogonal @3:00” experiments were used. All these configurations can be seen in Figure 6, where each box represents a different bearing configuration.
Figure 6: Bearing configurations used in the CWRU dataset. Each cell represents a specific acquisition setup containing two bearings—one on the drive-end and one on the fan-end. Fault types are denoted as follows: I for inner race fault, O for outer race fault, B for ball fault, and H for healthy.
In the division proposed by [5], faulty bearings with the same fault size are grouped, resulting in three different subsets (7, 14, 21 mils) that later are used to train and evaluate models. For example, when the subset with a fault size of 7 mils is used for training, the remaining subsets of 14 and 21 mils are used as test sets. This division prevents the occurrence of the same faulty bearings in both the train and test datasets, effectively addressing the data leakage issue. However, this solution is not entirely foolproof, as the dataset contains a single healthy bearing configuration, which must be split into the three previously described data subsets. In the [5] approach, this division is based on load, with healthy bearings at loads of 1, 2, and 3 HP being assigned to the 7, 14, and 21 mils subsets, respectively. This amounts to a condition-wise split of the healthy data, resulting in data leakage. Note that any division of the healthy bearing configuration on the CWRU dataset necessarily results in data leakage.
In contrast to the dataset division described above, in practice, faults may occur in various sizes, making it unrealistic to have a dataset distribution of only one fault size during training. This can lead to difficulty in training a model to detect faults at different sizes that did not appear in the training set. Therefore, it is important to consider a more diverse range of fault conditions in the dataset to ensure that the model can accurately predict them.
A more realistic data partitioning strategy involves creating subsets in which multiple loads, fault sizes, types, and locations appear randomly. In this context, we propose a split methodology that selects signals from a random fault size configuration for each fault location–type pair. To prevent data leakage, signals from the same bearing configuration (e.g., the same physical bearing under specific conditions) are restricted to either the training or testing set, but not both. The selected configurations form the test set, while all remaining signals are allocated to the training set.
Due to the limited number of healthy bearings in the dataset—only two, used alternately at the drive-end and fan-end depending on the faulty bearing location—our approach ensures that healthy data used for training and testing come from distinct locations. Specifically, the training set uses healthy signals from only one location, and the model is evaluated on the healthy signals from the opposite side.
In addition to this hold-out strategy, a 3-fold cross-validation procedure can be applied following the same selection principles, with the intent of using it to optimize hyperparameters. The first fold is generated by randomly selecting a fault size configuration for each fault location–type pair, as done in the hold-out scenario. The remaining two folds are then created by applying additional hold-out splits on the residual configurations, excluding the first fold. An illustration of this 3-fold setup is shown in Figure 7.
Figure 7: Illustration of the 3-fold split in the CWRU dataset following our proposed hyperparameter optimization methodology. Each cell contains two bearings: one on the drive-end and one on the fan-end.
While faulty bearings guide the primary structure of the splits, careful handling of the limited healthy data is critical. Given only two healthy bearings, each train-test configuration has two viable scenarios. In the first scenario, the healthy bearing on the fan-end is used in training, necessitating the exclusion of drive-end healthy signals from the training set since the evaluation will occur on that side. Notably, these excluded healthy signals cannot be reassigned to the test set, as they were acquired in sessions with a faulty bearing on the opposite end and may contain artifacts related to it. For the same reason, configurations involving both healthy bearings simultaneously are discarded. The second scenario is symmetrical: the model is trained using healthy signals from the drive-end and evaluated on the fan-end. This approach is illustrated in Figure 8.
Figure 8: Example of the two healthy-bearing split scenarios in our proposed methodology.
Using the 3-fold strategy shown in Figures 7 and 8, we generate three train-test combinations. Since each configuration involves training and testing on both sides (DE/FE), this results in six experiments in total. This methodology was adopted for model selection using the CWRU dataset. To perform evaluation, we created 50 additional 2:1 splits following the same principles illustrated in Figure 8. Each split supports two train-test configurations, yielding a total of 100 evaluation runs.
4.2.4 Summary
Table 3 summarizes the main characteristics of the bearing diagnosis datasets considered in this study. Note that, to facilitate the comparison of dataset diversity, we introduce the notion of [bearing, health state] instances. As discussed in the previous sections, the UORED-VAFCLS dataset associates each bearing with three distinct health states, two of which represent different fault severities. Consequently, although the dataset comprises 20 bearings, it yields 60 unique [bearing, health state] instances, offering greater diversity. In contrast, the CWRU and PU datasets provide a larger number of signals overall (specially PU), but only one health state per bearing, resulting in roughly three times fewer instances.
Table 3: Characteristics of the bearing diagnosis datasets.
Dataset Number of bearings Health states per bearing Number of [bearing, health state] instances Number of faulty instances Number of faulty instances in training
UORED-VAFCLS 20 3 60 40 24
CWRU 20333Excluding bearings with a 0.028” fault size. 1 18444We only consider instances with a faulty bearing accompanied by a healthy bearing at the opposite side (see Section 4.2.3). 18 12
Paderborn 17555These bearings correspond to the group with real faults, excluding those with combined faults. 1 17 11 7
Dataset Acquisition locations Conditions per instance Acquistions per condition per location Total number of signals
UORED-VAFCLS 1 1 1 60
CWRU 2 4 1 144
Paderborn 1 4 20 1360
4.3 Models and Training Details
4.3.1 Shallow Learning models with bearing fault based features
In shallow (i.e., non-deep) learning experiments, we adopt feature-based approaches using classical machine learning classifiers—specifically Random Forest and Support Vector Machines (SVM). The input features consist of well-established signal descriptors from the literature. Feature extraction is grounded in signal processing techniques applied to vibration data in both time and frequency domains, with prior studies demonstrating their effectiveness for bearing fault diagnosis.
In the work of [6], statistical metrics extracted from time-domain signals were employed, including mean, absolute median, standard deviation, skewness, kurtosis, crest factor, energy, RMS value, peak count, zero-crossing count, Shapiro-Wilk test statistic, and Kullback–Leibler divergence. In the frequency domain, [20] explored the use of bearing-specific fault frequencies, such as BPFI (Ball Pass Frequency of the Inner race), BPFO (Ball Pass Frequency of the Outer race), and BSF (Ball Spin Frequency). Both works reported favorable results using these handcrafted features, motivating their inclusion in our shallow learning pipeline. To highlight modulated fault components distributed across the spectrum, envelope analysis is performed to shift these components into the baseband, following the principles explained in [19]. Subsequently, algorithms are applied to extract the magnitudes of spectral peaks. In this study, the envelope spectrum is computed in the frequency range of 500 Hz to 10 kHz for the PU and UORED-VAFCLS datasets, which offer high sampling rates. For the CWRU dataset, limited to a 12 kHz sampling rate, a reduced band of 500 Hz to 6 kHz is adopted. Fault frequencies are computed based on the rotation speeds and geometric parameters provided within each dataset. Table 4 shows all features used on shallow learning models.
Table 4: Time and frequency domain hand-crafted features for Shallow Learning models.
Set Features
Time + Frequency
RMS, peak-to-peak, kurtosis, skewness, crest factor,
magnitude of all bearing fault frequencies (1x - 5x)
on envelope spectra
A notable limitation of shallow learning approaches lies in their reliance on metadata—particularly rotational speed (RPM) and bearing geometry—for the computation of specialized features. Furthermore, these models do not generalize naturally across datasets with differing fault types or label structures. As a result, we must train a separate classifier for each fault mode, increasing the complexity of model management. In contrast, deep learning models can learn representations directly from raw signals and are capable of handling varied datasets within a unified architecture.
Finally, our preprocessing of the training and test signals consisted of segmenting them into non-overlapping 1-second windows, resulting in 10 segments per signal, since all datasets contain 10-second acquisitions.
4.3.2 Deep Learning architectures
Informed by a review of contemporary deep learning models for vibration analysis, we selected the Wide First-layer Kernel Deep Convolutional Neural Network (WDCNN) [21] as the primary architecture for our investigation. This model, with approximately 172k parameters for an input segment length of 12,000 samples, offers a well-established baseline. To benchmark its performance against more recent or complex 1D convolutional architectures, a comparative analysis was conducted on the UORED-VAFCLS dataset, evaluating the WDCNN against the CDCN [22], WDTCNN [23], 1D-ConvNet [24], and RESNET1D [25] models. The results of this comparative study are presented in Section 5.1. To implement our multi-label classification framework, the output layer of each architecture was configured with a number of nodes equal to the number of detectable fault modes (excluding the healthy state). Each output node employs a sigmoid activation function, effectively operating as an independent binary classifier for a specific fault type.
A critical challenge posed by the CWRU and UORED-VAFCLS datasets is the limited number of acquisitions, which substantially increases the risk of model overfitting. To mitigate this, we implemented a data augmentation strategy combining two techniques. The first, Random Crop, generates training samples by extracting segments of a predefined length (e.g., 42,000 samples, corresponding to 1 second for UORED-VAFCLS) from random positions within the full signal. The second, Random Gain (referred to in [26] as Scaling), introduces stochastic amplitude variations by multiplying each signal by a scalar drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$. To ensure sufficient gradient updates per epoch, we created virtual datasets by applying these augmentations, effectively increasing the number of training examples per epoch by a factor of 20 for UORED-VAFCLS and 5 for CWRU. These techniques are applied exclusively to the training set, whereas the test set is preserved as non-overlapping segmented signals, as described in the previous section.
The augmentation hyperparameters were systematically optimized on the UORED-VAFCLS dataset. Following [26], the mean $\mu$ for the Random Gain distribution was fixed at 1.0, while the standard deviation $\sigma$ was tuned over the set 0.3, 0.5, 0.7. Our validation experiments revealed that the optimal configuration was a combination of Random Crop and Random Gain (RC+RG) with $\sigma=0.7$ for time-domain signals and $\sigma=0.3$ for frequency-domain inputs. To maintain a feasible computational budget, these optimized augmentation hyperparameters were subsequently applied without modification to all other datasets and experiments.
For the model selection phase (the inner loop of our CVM-CV protocol), we defined a hyperparameter search space to tune batch size, learning rate, and the input normalization strategy. Training was conducted for a fixed duration of 30 epochs across all experiments. The specific hyperparameter values explored during this tuning process are detailed in Table 5.
Table 5: Hyperparameter search space for model selection, applied across all datasets.
Batch size Learning Rate Normalization strategy
[16, 32, 64, 128, 256] [1e-2, 1e-3, 1e-4, 1e-5] [none, global, entry-wise]
Finally, three input representations were chosen: time-domain signals, frequency-domain (FFT) and envelope spectrum (FFT applied to temporal envelope). As mentioned in Section 4.3.1, the envelope spectrum highlights modulated fault components shifting them to baseband.
5 Experimental results
In this section, we apply the methodology described in Section 4 to each dataset. First, the best model architecture is identified through a CVM experiment on the UORED-VAFCLS dataset, comparing performances in the time and frequency domains. Next, we conduct evaluation experiments on all datasets, comparing three input representations across Deep Learning with two Shallow Learning models. Additionally, on the UORED-VAFCLS and PU datasets, we test different train–test split proportions to assess the impact of the number of bearings on a model’s generalization capacity.
5.1 UORED-VAFCLS dataset
The UORED-VAFCLS dataset served as the primary testbed to conduct a rigorous comparative analysis to identify the most effective deep learning architecture and input representation for bearing fault diagnosis. A comparative evaluation of five distinct 1D convolutional neural network architectures was performed using the CVM protocol on the 5 designated tuning splits. The objective was to determine the optimal combination of model, input representation (time-domain vs. frequency-domain), and hyperparameters. The results of this model selection phase, summarized in Table 6, reveal the leading performers.
Table 6: Best tuning results on the UORED-VAFCLS dataset across five architectures. Performance is reported as the mean and standard deviation of the Macro AUROC over the 5 designated tuning splits, along with the optimal hyperparameters identified.
Architecture Input Repr. Macro AUROC BS LR Normalization
1D-ConvNet Time 85.99% +- 7.80% 64 0.001 none
Frequency 81.25% +- 6.23% 256 0.0001 none
RESNET1D Time 76.05% +- 4.06% 32 0.01 none
Frequency 87.6% +- 10.13% 32 0.001 none
CDCN Time 82.45% +- 10.54% 16 0.001 none
Frequency 89.36% +- 10.12% 32 0.0001 none
WDTCNN Time 91.09% +- 5.99% 32 0.01 none
Frequency 87.32% +- 5.12% 256 0.01 none
WDCNN Time 91.47% +- 4.87% 128 0.001 none
Frequency 93.24% +- 5.93% 256 0.0001 none
The WDCNN architecture demonstrated superior performance, achieving the highest mean Macro AUROC scores for both time-domain (91.47%) and frequency-domain (93.24%) inputs. The WDTCNN also yielded strong results with time-domain inputs (91.09%). Based on its consistently high performance and robustness across both input modalities, the WDCNN was selected as the primary architecture for the full performance evaluation.
The final evaluation was conducted using the selected WDCNN model on the 100 disjoint test splits. The results, presented in Table 7, confirm the superiority of the frequency-domain representation, which achieved a Macro AUROC of 93.12% ± 4.26%. This result serves as the primary deep learning benchmark for this dataset. In parallel, we also report the performance of shallow learning models trained on handcrafted features.
Table 7: Final evaluation results on the UORED-VAFCLS dataset for the selected WDCNN model and corresponding shallow learning benchmarks. Performance is the mean and standard deviation of the Macro AUROC over 100 disjoint test splits.
Model Input repr. / Features Macro AUROC
WDCNN Time 90.69% ± 5.43%
Frequency 93.12% ± 4.26%
Envelope Spectrum 89.57% ± 5.34%
Random Forest Time + Frequency 85.58% ± 4.74%
SVM Time + Frequency 81.33% ± 8.61%
A central tenet of our methodology is the critical role of the number of unique bearings in the training set for model generalization. To empirically validate this principle, we conducted a sensitivity analysis on the UORED-VAFCLS dataset by systematically varying the train-to-test bearing ratio, exploring configurations of 1:4, 2:3, 3:2, and 4:1. To isolate the effect of bearing diversity, it was critical to control the total volume of training data seen by each model. We achieved this by varying the number of steps per epoch for each split configuration, ensuring that all models processed approximately the same number of samples used in the baseline experiment (3:2) during training and applying the same Data Augmentation techniques. This methodology disentangles the influence of data diversity from the confounding variable of data quantity, clarifying whether performance changes are due to a richer dataset or simply a larger one.
Figure 9: Impact of train-test split ratio on model performance on the UORED-VAFCLS dataset. The plot shows the mean Macro AUROC (and standard deviation as error bars) calculated over 100 evaluation splits for four distinct bearing-level train-to-test ratios.
As hypothesized, a clear trend emerged. As illustrated in Figure 9, increasing the number of training bearings (e.g., a 4:1 ratio) leads to higher mean Macro AUROC scores, as the model benefits from greater component-level diversity during training. However, this comes at the cost of a smaller and less diverse test set, which may limit the reliability of the evaluation—especially in real-world scenarios where generalization to unseen components is critical. On the other hand, splits with more bearings in the test set (e.g., 1:4) provide a broader basis for evaluation, but may yield lower performance due to the reduced diversity in training data. These findings highlight the importance of selecting a partitioning strategy that not only supports model generalization but also ensures that performance estimates are based on sufficiently diverse and representative test sets.
5.2 PU dataset
The experiments conducted on the curated Paderborn University (PU) dataset revealed a significant generalization challenge, with model performance approaching that of a random classifier in both the time and frequency domains. Both input representations exhibited rapid overfitting, which necessitated adding a Dropout layer with a 90% dropout rate before the output layer to regularize the model.
The final evaluation results, presented in Table 8 quantify the extent of the generalization challenge. The low mean scores and high standard deviations underscore the model’s inability to consistently learn a robust fault signature from the limited training data.
Table 8: Evaluation results on the curated PU dataset. Performance is reported as the mean and standard deviation of the Macro AUROC over 100 disjoint test splits.
Model Input Repr. / Features Macro AUROC
WDCNN (with Dropout) Time 59.51% ± 14.14%
Frequency 61.89% ± 13.65%
Envelope Spectrum 80.16% ± 14.48%
Random Forest Time + Frequency 69.67% ± 15.73%
SVM Time + Frequency 64.42% ± 15.83%
Our primary hypothesis for this poor performance is the limited component diversity in the curated training subset of naturally degraded bearings, especially when compared with the broader variability of the UORED-VAFCLS dataset (Table 3). We first considered the possibility of model underfitting (i.e., insufficient model capacity), but this was ruled out, as the WDCNN consistently achieved near-perfect performance on the training partitions, demonstrating its ability to fit the available data. This observation leads us to conclude that the issue lies not with the model, but with the data itself.
While each bearing was recorded under four different operating conditions, our results suggest that this intra-bearing variation is insufficient to compensate for the limited inter-bearing diversity, which is manifested mostly in splits where the model performances reach values around 0.5. In these cases, the unique physical signature of each bearing appears to be the dominant feature, making it difficult for the model to generalize to unseen components. On the other hand, results with Envelope Spectrum, a more specialized input representation, showed us that specialized features may contribute to reduce the impact of memorization.
To empirically validate this diversity-limitation hypothesis, we replicated the split-proportion analysis performed on the UORED-VAFCLS dataset. As depicted in Figure 10, a clear and monotonic increase in the mean Macro AUROC was observed as the proportion of bearings in the training set increased. Although the overall performance remains low, this direct correlation provides strong evidence that the lack of component diversity is the principal bottleneck limiting model generalization on this dataset.
Figure 10: Impact of train-test split ratio on model performance on the PU dataset using envelope spectrum as the input representation.
5.3 CWRU dataset
The results obtained on the CWRU dataset were comparable to those observed on the PU dataset for deep learning approaches, with the envelope spectrum domain yielding the best overall performance. In contrast, shallow learning models—particularly the Random Forest—achieved better results, as shown in Table 9.
Table 9: Macro AUROC (mean ± standard deviation) results obtained applying our methodology in the CWRU dataset using Deep and Shallow Learning Models.
Model Input Repr. / Features Macro AUROC
WDCNN Time 63.22% ± 10.24%
Frequency 70.90% ± 8.95%
Envelope Spectrum 74.51 ± 9.43%
Random Forest Time + Frequency 84.40% ± 9.31%
SVM Time + Frequency 76.42% ± 9.87%
We hypothesize that the limited performance of deep learning models may stem from their tendency to memorize specific bearing signatures seen during training, rather than generalizing to unseen signals. This effect is likely amplified in datasets with limited variability, such as CWRU. We speculate that in more diverse datasets, the deep models’ feature extractors (backbones) would be exposed to a broader range of examples, enabling the learning of more general and predictive representations. In contrast, for the CWRU dataset, the use of handcrafted features—both specialized and general—proves to be a more effective strategy.
Finally, we compare two distinct train-test splits on CWRU (1:2 and 2:1), following the same procedure as before, seeking to observe the impact of different split ratios on model performance. For this experiment, in order to guarantee an equal number of samples in both splits, we applied segmentation with a $12000$ size and an overlap of $53\%$, totalizing $20$ segments per signal in the 1:2 split. Figure 11 shows the results of this experiment which, although limited due to available number of configurations, is in line with those obtained for PU and UORED-VAFCLS, suggesting that increased diversity can be beneficial to performance.
Figure 11: Impact of train-test split ratio on model performance on the CWRU dataset using Random Forest with hand-crafted features.
6 Experiments with data leakage
In this section, we investigate the effects of data leakage across multiple datasets by designing experiments that fit their unique structures. Earlier, we showed using a toy problem that data leakage produces overoptimistic results, which are not suitable for decision-making. We now present similar findings using real data.
Although previous studies have investigated data leakage in bearing diagnosis, to the best of our knowledge, no prior work has done so solely by altering the test set. In our proposed experiments, we address this by keeping the training set—and thus the model—fixed, while modifying only the test set. This is important since changing both sets introduces a confounding factor: it becomes impossible to know whether a given result is inferior because leakage was avoided or because training diversity was reduced. We eliminate this confounding factor with our setup, in which the presence or absence of data leakage is determined entirely by the test set composition. This approach also explains why our no-leakage results differ from those presented in Section 5, as we designed specific experiments for each dataset according to its inherent limitations.
Our experimental design for all datasets contrasts a leakage-free evaluation protocol with common but flawed partitioning methods that introduces data leakage.
6.1 UORED-VAFCLS dataset
The experimental protocol for the UORED-VAFCLS dataset was structured around 10 distinct train-test splits, each following our standard partitioning strategy presented in Section 4.2.1. In every split, the training set remained fixed across all test scenarios to ensure a controlled comparison. Specifically, training was performed using only the first 80% of each signal from the training bearings. Additionally, to simulate a realistic data leakage scenario, all “severe fault" signals from train bearings were removed from the original train set and set aside for use in one of the leakage scenarios.
Based on this fixed training set, we evaluated two distinct leakage scenarios:
  • Segmentation-level leakage: This scenario examines leakage within individual signals. While the training set only included the first 80% of each signal from the training bearings, the test set in this condition was constructed exclusively from the remaining 20% of those same signals—thus exposing the model to future portions of time series it had already partially seen during training.
  • Bearing-level leakage: In this scenario, there is no segmentation-level leakage, but the test set includes signals from bearings that also appear in training. To create this, the previously removed “severe fault” segments from train bearings were reintroduced into the test set. To maintain the test set size and class balance, an equal number of severe fault test samples were removed and replaced by these leakage samples.
Table 10: Performance comparison on the UORED-VAFCLS dataset, demonstrating the impact of bearing-level and segmentation-level data leakage versus the proposed leakage-free methodology.
Split Model Input Repr. / Features Macro AUROC
No leakage WDCNN
Time
Frequency
91.80% ± 5.22%
74.81% ± 11.54%
Bearing-level leakage WDCNN
Time
Frequency
94.18% ± 3.50%
81.22% ± 6.22%
Segmentation-level leakage WDCNN
Time
Frequency
99.15% ± 7.19%
90.11% ± 5.42%
The results, summarized in Table 10, demonstrate the substantial and misleading performance inflation caused by both forms of data leakage. Introducing the bearing-level leakage resulted in an artificial performance increase of approximately 2.4% for time-domain inputs and 6.4% for frequency-domain inputs. The more severe segmentation-level leakage produced an higher performance increase, elevating the Macro AUROC by 7.3% and 15.3% for time and frequency representations, respectively. These findings provide compelling empirical evidence for the necessity of strict, bearing-wise data partitioning to ensure the validity and reliability of model evaluation in bearing fault diagnosis.
6.2 PU dataset
For the experimental evaluation on the PU dataset, we also generated 10 distinct data splits. In each split, the training set was constructed utilizing data from three of the four available operating conditions. From these selected conditions, 15 measurement repetitions were allocated for model training, while the residual 5 repetitions, along with all data from the fourth operating condition, were reserved to be used in the leakage scenarios. As a key modification, we truncated all training signals by removing the final 25% of each recording. Using this setup, we proceeded to test our model against three types of data leakage. First, to induce segmentation-level leakage, we followed a procedure similar to that used for the UORED-VAFCLS dataset, but now placing the 25% segments taken out of the training set into the test set (instead of 20%). The remaining types of data leakage are detailed below:
  • Bearing-level leakage (Condition-wise split): In this scenario, the model is exposed to the same physical bearing in both the training and testing sets, but under different operating conditions. Each training set has three of the four available conditions, while the remaining one is added to the test set, introducing leakage.
  • Bearing-level leakage (Repetition-wise split): This represents a more severe leakage case. Here, the model is trained on one measurement repetition and tested on a different repetition from the exact same bearing and operating condition. We simulate this type of leakage by reserving 15 of the 20 measurement repetitions for model training, while adding the rest to the test set.
Table 11: Performance comparison on the PU dataset, demonstrating the impact of bearing-level and segmentation-level data leakage versus the proposed leakage-free methodology.
Split Model Input Repr. / Features Macro AUROC
No leakage WDCNN
Time
Frequency
53.2% ± 17.2%
53.3% ± 16.5%
Bearing-level leakage (Condition) WDCNN
Time
Frequency
85.9% ± 14.9%
92.4% ± 11.3%
Bearing-level leakage (Repetition) WDCNN
Time
Frequency
98.3% ± 2.6%
99.9% ± 0.1%
Segmentation-level leakage WDCNN
Time
Frequency
98.2% ± 3.2%
100.0% ± 0.0%
The results in Table 11 reveal a similar pattern with the UORED-VAFCLS experiments, showing the impact of bearing-level leakage (with different partitioning methods) and segmentation-level leakage. Particularly, we observe the key difference between leakage across condition- and repetition-wise splits. The provided results shows that repetitions of the same experiment provide an easy way to memorize signal characteristics (a gain of 46.6% on frequency domain), sharing similar results with segmentation-level leakage, that reached perfect performance. Meanwhile, the condition-wise split also shows a high performance gain of 39.1% compared to the valid experiment, indicating that this type of data splitting also provides an easy way for model memorization.
6.3 CWRU dataset
The experimental protocol for the CWRU dataset consists in training with samples grouped by condition and fault size, as represented in Figure 12. In total, there are 12 groups, where every experiment was performed training with a single group and testing on the others. In our leakage-free protocol, the models were evaluated in groups with a different fault size, which corresponds to different bearings (e.g A $\rightarrow$ E). With bearing-level leakage, the test set consisted of the three remaining groups on each line (e.g A $\rightarrow$ B, A $\rightarrow$ C and A $\rightarrow$ D). Finally, with segmentation-level leakage, we repeated the same process as in UORED-VAFCLS, where all models were trained with only 80% of the group signals, leaving 20% separated to be used in the test set.
Figure 12: CWRU leakage experiment groups.
The results are presented in Table 12 and follow the same pattern as the UORED-VAFCLS and PU datasets. In this case, both scenarios of bearing-level leakage (with condition-wise splits) and segmentation-level leakage resulted in perfect or near-perfect scores with both time and frequency domain inputs. This reinforces our hypothesis that condition-wise splits induces a strong data leakage, very similar to segmentation-level leakage on the CWRU dataset.
Table 12: Performance comparison on the CWRU dataset, demonstrating the impact of bearing-level and segmentation-level data leakage versus the proposed leakage-free methodology.
Split Model Input Repr. / Features Macro AUROC
No leakage WDCNN
Time
Frequency
66.4% ± 16.8%
62.4% ± 20.6%
Bearing-level leakage (Condition) WDCNN
Time
Frequency
99.9% ± 0.2%
100.0% ± 0.0%
Segmentation-level leakage WDCNN
Time
Frequency
99.8% ± 0.2%
99.9% ± 0.1%
7 Discussion and conclusion
In this work, we demonstrated that performance evaluations in bearing fault diagnosis can be significantly overestimated when data leakage is present. We applied a rigorous and leakage-free methodology to three widely used datasets—CWRU, PU, and UORED-VAFCLS—and observed substantial differences in model performance depending on dataset diversity, input representation, and learning approach.
Our experimental results demonstrate that when using a leakage-free methodology, deep learning models do not exclusively yield the best performance. In fact, shallow models trained with handcrafted features can be highly competitive and, in some scenarios, even superior. This is clearly illustrated by our findings on the CWRU dataset, where a Random Forest model achieved a Macro AUROC of 85.52% ± 7.43%, outperforming all deep learning baselines. Even when not the top performer, such traditional techniques provide a crucial performance baseline. Conversely, on the PU dataset, the WDCNN with spectrum envelope input yielded the best performance (80.16% ± 14.48%), highlighting the importance of choosing a good input representation. Finnaly, on the UORED-VAFCLS dataset, WDCNN with frequency-domain input outperformed all other configurations, reaching a Macro AUROC of 93.12% ± 4.26%.
These findings suggest that model performance is highly dependent on both the characteristics of the dataset and the choice of representation and architecture. For example, PU showed lower and more variable results, likely due to limited variability in bearing types. In contrast, UORED’s higher scores suggest it may be a better benchmark for assessing model generalization.
Our methodology also highlights the difficulty of avoiding data leakage in the CWRU dataset. Despite being the most widely used benchmark in bearing fault diagnosis, CWRU presents significant challenges for rigorous evaluation due to its structure. In contrast, the PU and UORED-VAFCLS datasets are more amenable to clean bearing-wise splits, making them more suitable for assessing true model generalization.
Based on these insights, we recommend the following to researchers in the field:
  • Bearing-wise split: Ensure strict separation of bearings across training and test splits to avoid misleading, inflated performance estimates due to data leakage.
  • Datasets with different properties: We encourage the community to systematically test models on datasets with varying properties (e.g., PU, UORED-VAFCLS) to better assess robustness, as well as to explore multiple diversity configurations within each dataset. It is also crucial to select datasets with a structure that supports proper bearing-wise splitting and contain a sufficient number of bearings per class to enable meaningful evaluation.
  • Model selection and evaluation protocol: The process for tuning hyperparameters must be clearly detailed. These parameters must be chosen without using the evaluation performance.
  • Multi-label formulation: This formulation provides a setup for diagnosing co-occurring faults, a realistic scenario, and enables a more precise evaluation with prevalence-independent metrics, such as the Macro AUROC metric.
  • Models: We suggest that a comprehensive evaluation should not default to deep learning architectures. Shallow models often provide competitive or even superior results depending on the dataset, and we encourage their inclusion as robust baselines.
  • Share code, data splits, and evaluation pipelines: Reproducibility is key for advancing the field and ensuring fair comparisons between proposed methods.
By following these practices, the community can foster more reliable and generalizable machine learning systems for bearing fault diagnosis applications.
[…]
Acknowledgements
This research was supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), grants 164299/2021-1 and 304619/2022-1, and the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES), grants 88887.951193/2024-00 and 88887.137796/2025-00.
Appendix A Derivation of the maximum achievable accuracy in the toy example
Consider a discrete random variable $y\in\{0,1\}$ representing the state of a given bearing, where $y=0$ denotes the healthy state and $y=1$ denotes the faulty state.
We observe a continuous random vector $X\in\mathbb{R}^{N}$ representing $N$ fault predictive features, whose componentes are i.i.d. Gaussian random variables conditioned on $y$, given by $X_{i}\mid y\sim\mathcal{N}(yA,1)$, i=1,2,,Ni=1,2,...,N. More precisely,
$P({X}\mid y=0)=\prod_{i=1}^{N}\sqrt{\frac{1}{2\pi}}e^{-\frac{{X_{i}}^{2}}{2}}\quad\text{and}\quad P({X}\mid y=1)=\prod_{i=1}^{N}\sqrt{\frac{1}{2\pi}}e^{-\frac{({X_{i}}-A)^{2}}{2}}.$ (1)
It is well-known that the maximum a posteriori (MAP) decision rule minimizes the error probability (and therefore maximizes accuracy). The MAP rule selects the hypothesis $\hat{y}$ for the bearing health state as $\hat{y}=\arg\max_{y}P(y\mid X)$. In this case, this amounts to deciding $\hat{y}=1$ whenever $P(y=1\mid{X})>P(y=0\mid{X})$. It follows that
$\hat{y}=1$ $\iff P(y=1\mid{X})>P(y=0\mid{X})$ (2)
$\iff\frac{P({X}\mid y=1)P(y=1)}{P({X})}>\frac{P({X}\mid y=0)P(y=0)}{P({X})}$ (3)
$\iff\frac{P({X}\mid y=1)}{P({X}\mid y=0)}>\frac{P(y=0)}{P(y=1)}$ (4)
$\iff e^{\sum_{i=1}^{N}\frac{-(X_{i}-A)^{2}}{2}+\frac{X_{i}^{2}}{2}}>\frac{P(y=0)}{P(y=1)}$ (5)
$\iff\sum_{i=1}^{N}\frac{-(X_{i}-A)^{2}}{2}+\frac{X_{i}^{2}}{2}>\log\frac{P(y=0)}{P(y=1)}$ (6)
$\iff A\sum_{i=1}^{N}X_{i}-\frac{NA^{2}}{2}>\log\frac{P(y=0)}{P(y=1)}$ (7)
$\iff\frac{1}{N}\sum_{i=1}^{N}X_{i}>\frac{1}{NA}\log\frac{P(y=0)}{P(y=1)}+\frac{A}{2}.$ (8)
Since we assume $P(y=0)=P(y=1)=\frac{1}{2}$, we have that, under the MAP rule,
$\hat{y}=\begin{cases}0,&\text{if }\bar{X}\leq A/2,\\ 1,&\text{if }\bar{X}>A/2,\end{cases}\quad\text{where $\bar{X}=\frac{1}{N}\sum_{i=1}^{N}X_{i}$}.$ (9)
We now wish to find the corresponding accuracy $P(\hat{y}=y)$. Note that $\bar{X}\mid y\sim\mathcal{N}(yA,1/N)$. Since
$P(\hat{y}=0\mid y=0)$ $=P(\bar{X}\leq A/2\mid y=0)=\Phi\left(\frac{A/2}{1/\sqrt{N}}\right)=\Phi\left(\frac{A\sqrt{N}}{2}\right)$ (10)
$P(\hat{y}=1\mid y=1)$ $=P(\bar{X}>A/2\mid y=1)=1-\Phi\left(\frac{A/2-A}{1/\sqrt{N}}\right)=1-\Phi\left(\frac{-A\sqrt{N}}{2}\right)=\Phi\left(\frac{A\sqrt{N}}{2}\right)$ (11)
where $\Phi$ denotes the standard Gaussian c.d.f., it follows that
$P(\hat{y}=y)=P(\hat{y}=0\mid y=0)P(y=0)+P(\hat{y}=1\mid y=1)P(y=1)=\Phi\left(\frac{A\sqrt{N}}{2}\right).$ (12)
For the special case of $N=3$ and $A=1.5$, we have $P(\hat{y}=y)=\Phi\left(\frac{1.5\sqrt{3}}{2}\right)\approx 0.9030$.